# A simple model of globally magnetized accretion discs

### v1

@@ 2.4 Resistive and viscous dissipation
There are two sources of dissipation in our model: viscous and resistive. The viscous dissipation rate per unit area is simply given by
@@
To calculate the resistive dissipation, we evaluate
@@
expresses the dissipation of $B_{\phi}$ and the other term refers to $B_{r}$. In discs supported vertically by non-magnetic pressure, as in standard thin discs (with $q\sim O(1)$), $g\gg 1$ and the dissipation due to $B_{\phi}$ dominates. In this limit, with $\xi\sim O(1)$, the resistive and viscous dissipation rates are comparable. In a steady state, we have
@@
$F_{\rm acc}={1\over 2}\rho rH\bar{v}\Omega^{2}={\xi\over 2}\eta\rho r\Omega^{2}\ ,$ (29)
where we have used the accretion speed in the steady-state limit. As noted by Ferreira & Pelletier (1993), the ratio of dissipation to the local release of binding energy is $\sim\xi^{-1}(H/r)$, i.e., small whenever the inflow speed is strongly boosted by magnetic torques.
3 MRI Turbulence Closure
We have seen that the evolution of our simple slab disc model depends critically on the strength of turbulence expressed through the parameter $q$, which represents the ratio of $v_{{\rm A}z}H$ to the turbulent resistivity $\eta$ or, in physical terms, the ratio of the resistive diffusion time to the Alfvén crossing time. While the complete dependence of the model on $q$ is complex (and probably reflects the model’s limitations), for $q\ga 1$ we see that increasing $q$ (weaker turbulence) leads to stronger advection and retention of magnetic flux, through both faster inflow speeds (due to magnetic extraction of angular momentum) and weaker resistive diffusion.
Moreover, these behaviors can be very sensitive to the level of turbulence when $q\approx\pi/2$. While the numerical values at which the behaviors set in are undoubtedly an artifact of the simplified model, the qualitative trends are likely robust and can be understood in terms of the slightly different behaviors of mass advection and magnetic flux advection in a resistive disc, and their delicate balance.
@@ 3 MRI Turbulence Closure
[…]
However, the simulations also indicate that the MRI-driven turbulent energy density approaches and exceeds the gas pressure when the vertical magnetic field is still quite weak, at $\beta_{z}\la 10^{2}$. The enhanced turbulent pressure will thicken the disc (Scepi et al., 2024), but also — according to our disc dynamo model — drive the Alfvén speed associated with the organized toroidal field, $v_{{\rm A}\phi}$, above the sound speed. In this limit, the growth rate of MRI is suppressed (Blaes & Balbus, 1994). Quasi-linear arguments suggest that the saturated level of MRI-driven turbulence is proportional to the instability growth rate on scales $\sim H$ (Begelman & Armitage, 2023), in which case the amplification of $B_{\phi}$ as $\beta_{z}$ decreases could drive $q$ towards or even beyond $\pi/2$, leading to a steady state (if $q\la\pi/2$) or the rapid accumulation of flux (if $q>\pi/2$). We will adopt the ansatz that for weakly magnetized discs, there is a critical $\beta_{z}$, $\beta_{\rm crit}\sim 10^{2}$, at which $v_{B}=0$ and a steady state can be maintained for both mass flux and magnetic flux.
@@ 4.1 Weakly magnetized discs
Adopting the closure relation for this case, we take $\beta_{z}=\beta_{\rm crit}=10^{2}\beta_{2}$, implying
@@
This is also very close to the mean accretion speed $|\bar{v}|$, from equation (23). We assume that the disc is supported vertically by gas pressure, implying $H=c_{s}/\Omega$. Interestingly, the reduced level of dissipation inside the disk implies that a magnetically boosted, steady state disc cannot be supported by radiation pressure, as can be seen by the following argument. Vertical support of a disc by radiation pressure requires a local flux $F_{d}(r)=(H/r)GMc/\kappa r^{2}$, where $\kappa$ is the opacity. But we have seen in §2.4 that the actual disc flux is roughly a fraction $H/r$ of the locally liberated binding energy, or $\sim(H/r)GM\dot{M}/8\pi r^{3}$. Comparing the two fluxes, we find that maintaining radiation pressure support requires $\dot{M}$ to be a specific function of radius, $\dot{M}\sim 8\pi cr/\kappa$. This condition clearly cannot be satisfied at all $r$ in a conservative accretion flow, with constant $\dot{M}$. [^1] Expressed as $r\sim\kappa\dot{M}/8\pi c$, this is the condition for the “trapping radius,” (Begelman, 1978), where advection and diffusion of radiation are comparable.
The accretion rate, given by $\dot{M}=4\pi\rho rH|\bar{v}|$, can be normalized to the Eddington value using the electron scattering opacity, $\kappa_{\rm es}$:
@@
The dissipation is the sum of resistive and viscous contributions, and is given (for $\xi={\rm Pr_{t}}=1$) by
@@
These scalings (and others that can be derived from them) resemble the standard accretion disc scalings in the scattering-dominated regime. In both opacity limits, however, the systematic deviation of the results from standard disc models embody the lower radiative efficiencies of magnetized discs and the suppression of radiation pressure as a major source of vertical support, when $\dot{m}<1$.
@@ 4.2 Strongly magnetized “magnetically elevated” discs
5 Stability
While a full stability analysis of our disc evolution equations is beyond the scope of this paper, we can examine the local stability of our steady-state models by perturbing them with short-wavelength radial perturbations of the form $y=y_{0}(1+\delta y)$ with $\delta y\propto e^{i(kr+\omega t)}$ and $H\ll k^{-1}\ll r$. We will take $\delta\xi=0$, which is plausible since the external field configuration (and hence the inclination of the poloidal field at the disc surface) is probably governed by global external boundary conditions. Defining $\nu\equiv(\omega/k)(H_{0}/\eta_{0})$, we write the perturbed mass and flux equations in the form
$\nu(\delta\rho+\delta H)=\delta\rho+\delta H+\delta\bar{v}\ ,$ (51)
$\nu\delta B_{z}=-\delta v_{B}\ ,$ (52)
$\nu\left(\delta v_{{\rm A}z}+{1\over 2}\delta\rho\right)=\left({\pi\over 2}\right)^{2}\delta q\ ,$ (55)
5.1 Weakly magnetized discs
5.2 Strongly magnetized discs
A self consistency condition on this stable regime leads to a lower limit on disc thickness. For $c_{s}\la v_{{\rm A}\phi}\la(c_{s}v_{k})^{1/2}$, the maximum MRI growth rate decreases with $v_{{\rm A}\phi}$ (Blaes & Balbus, 1994), actually vanishing for $v_{{\rm A}\phi}=2\sqrt{2}(c_{s}v_{K})^{1/2}$ when $B_{\phi}\propto r^{-5/4}$ (Pessah & Psaltis, 2005; Das et al., 2018). According to the saturation argument proposed by Begelman & Armitage (2023), $\eta\propto v_{{\rm A}\phi}$ in this regime (provided that MRI is responsible for the turbulence), implying $b\sim-1$ and instability. In order to enter the stable regime of disc evolution, $H$ must satisfy
@@ 6 Astrophysical Context
A premise of this paper is that accretion discs could be supplied with net magnetic flux from their surroundings. We have argued that the ability of a disk to concentrate flux by advecting it inward is sensitive to its level of magnetization (e.g., through the plasma $\beta$-parameter). In this section we address the global question of whether a disc forced into a highly magnetized state at large radii will then evolve toward a similar state at smaller radii. Given the unique radial dependence of the magnetic field, $B_{z}\propto r^{-5/4}$, required for steady accretion in a strongly magnetized disc, this question is clearly crucial for assessing the ability of discs to enter the MAD state.
@@ 6.1 Shakura-Sunyaev–like discs
If $\dot{M}$ is so low that either of the conditions (60) or (61) is satisfied, then the standard Shakura-Sunyaev mode should not be self-consistent because magnetic stresses will take over both the vertical and angular momentum transport. For accretion rates between the two limiting values, we would expect vigorous turbulence with dynamo effects leading to a magnetically supported disc with a magnetically boosted inflow speed, both of which would exacerbate the departure from viscous disk behavior. Paradoxically, in the more extreme limit of condition (61), we face the opposite situation that the level of turbulence might drop significantly as instability is quenched by the strong vertical field; in this case, the disc might effectively stop accreting.
@@ 6.2 Formation of MAD-like states
Like the solution with $\dot{M}_{0}>\dot{M}_{\rm mag}$, this steady flux advection solution also cannot persist to arbitrarily small radii, since the disc scale height satisfies $H/r\propto vr^{1/2}\propto r^{3/4}$. As before, we would expect the strong magnetization assumption to fail when $v\la 0.1c_{s}$, but unlike the case with high $\dot{M}_{0}$, the low-$\dot{M}_{0}$ state cannot revert to a viscous disk solution since strong flux advection continues to be maintained from the outer boundary. Instead, we expect flux to continue to accumulate, working its way inward until the flow approaches a steady state with $\dot{M}\sim\dot{M}_{0}$ and MAD scaling all the way to the central object. Note that the eventual state of the disc would the true steady state, with $v_{B}\approx 0$, as opposed to the steady flux-advection cases we have been considering.
This picture thus provides a possible scenario for creating a steady-state MAD. However, it assumes that the toroidal field is able to self-regulate to $B_{\phi}\sim v_{K}\rho^{1/2}<B_{z}$. There is no guarantee that this will happen, and we should consider the alternate possibility that the disc “freezes” near $r_{0}$, with so little dynamo activity at these high magnetization levels that $B_{0}$ is inadequate to drive accretion at even the relatively modest rate of $\dot{M}_{0}$. In this case, we would expect the column density near $r_{0}$ to build up until it reaches a level where strong dynamo activity resumes, $B_{\phi}$ approaches $\sim 10B_{z}$, and a period of MAD-like accretion with $\dot{M}\sim\dot{M}_{\rm mag}$ occurs. Since the mass supply rate is smaller than $\dot{M}_{\rm mag}$, this would have to be episodic accretion with an on-off duty cycle of $\dot{M}_{0}/\dot{M}_{\rm mag}$.
@@ 7 Discussion and conclusions
We have presented a simple, analytic dynamo model for thin accretion discs containing a unidirectional vertical magnetic field, $B_{z}(r)$. Because these discs are threaded by indestructible net magnetic flux, mean-field stochastic dynamo effects are not essential for maintaining the field and we ignore all of the commonly studied terms with the exception of an isotropic turbulent resistivity, $\eta$, which is related to the turbulent viscosity via a turbulent magnetic Prandtl number, ${\rm Pr}_{t}$. Importantly, we make no claims whatsoever as to the actual role of mean-field or other small-scale dynamo effects; they could well overwhelm the effects we discuss here. However, if they do turn out to be relatively unimportant, our analysis shows that the large-scale field structure can be robustly governed by the balance between field amplification by shear and radial flow, versus resistive diffusion and dissipation.
To isolate the basic features of the dynamo, we consider a local slab geometry with a scale height $H(r)$, with density $\rho(r)$ and $\eta(r)$ constant for $-H<z<H$ and zero elsewhere. Assuming fixed angular velocity $\Omega(r)$ (taken to be Keplerian in our applications), we calculate the vertical dependences of $B_{r}$, $B_{\phi}$ and the radial velocity explicitly from the induction and angular momentum conservation equations. Both $B_{r}$ and $B_{\phi}$ exhibit sinusoidal behavior about the midplane, with $B_{\phi}$ showing additional linear behavior with $z$ due to the viscous stress and advection of radial magnetic flux.
@@
The sensitivity of the model behavior to $q$ underscores the importance of understanding the relationship between small-scale instabilities and large-scale magnetic structures. Our lack of such understanding is most acute in the strongly magnetized cases. For the closure we have used in this limit — which assumes that the maximum pressure of the dynamo field is comparable to the turbulent energy density — the squeezing effect of the dynamo field close to the midplane is of order unity, and is thus quantitatively but not qualitatively important. This begs the question of that happens if $B_{\phi}^{\prime}$ reverses sign within the disc, or indeed how closely the level of turbulence is linked to the local behavior of the dynamo field. Our closure seems reasonable on energetic grounds, since the same shear flow is responsible for both the turbulent and organized field. But the argument from which the closure relation is derived (Begelman & Armitage, 2023) addresses a more specific context in which the turbulence in a highly magnetized disc is not driven directly by shear via MRI (which Begelman & Armitage 2023 argues becomes too weak when the disc is highly magnetized), but rather by tearing instabilities driven by the vertical gradient of $B_{\phi}$, which in turn has been set up by the shear. Since the strength of the instability depends on $B_{\phi}^{\prime}$, might not the strength of the turbulence and thus $\eta$ be tied to the field structure locally, with dramatic variations as a function of height?
@@ 7.1 Midplane outflows and field reversals
The trigonometric behavior of field strengths and radial velocities with height in our dynamo model suggest that discs may exhibit internal radial and toroidal field reversals, accompanied by zones of outward radial flow. Several simulations of magnetized have exhibited accretion dominated by regions offset from the midplane (Zhu & Stone, 2018; Lančová et al., 2019; Mishra et al., 2020), accompanied by weak inflow, or even outflow (“backflow”) near the equator. Analyzing the effect in their simulations, Zhu & Stone (2018) argue that this accretion pattern is dominated by the $\phi z$ stress due to large-scale magnetic fields, as earlier noted by Beckwith et al. (2009), rather than by viscosity operating on vertical shear (Kluzniak & Kita, 2000; Guilet & Ogilvie, 2012, 2013). For our magnetically boosted models (i.e. neglecting internal viscosity), we find midplane outflow for $(2n-1)\pi<q<2n\pi$, where $n\geq 1$ is an integer. The corresponding disc exhibits $n$ internal field reversals. Such high values of $q$ would correspond to depressed levels of turbulent resistivity compared to the fiducial scale $v_{{\rm A}z}H$; it remains to be seen under what conditions this would apply. We note that the source of angular momentum (and energy) driving midplane outflow is the accretion occurring further from the midplane. A portion of the angular momentum liberated in the inflow zone is deposited in the midplane region, driving outflow, while the remainder is carried away by magnetic stresses at the disc surface. In all cases, the vertically integrated mass flux is inward and the layer nearest the disc surface is accreting.
The direction of flow can also be associated with the geometry of the poloidal field. Changing the sign of $v_{r}(0)$ changes the sign of both $B_{r}$ and $B_{\phi}$ near the midplane. Given our magnetic sign convention $B_{z}>0$, equatorial outflow corresponds to $B_{r}<0$, which means that the poloidal field lines projected in the azimuthal plane should bend inward away from the midplane. This is clearly seen in Figure 3 of Mishra et al. (2020), where the field lines bend inward in a region around the midplane characterized by weak outflow. Moreover, our model predicts that $|B_{r}|$ should be maximized where $v_{r}$ changes sign and vanish in the middle of an elevated accretion zone — both effects also apparent in Mishra et al. (2020) Figure 3. This behavior gives the poloidal field lines the appearance of being pinched inward at some height above and below the midplane. Analogous information is displayed in Figure 15 of Zhu & Stone (2018) where pinching is also seen, although the relationship between $B_{r}$ and $v_{r}$ is not so clear. Zhu & Stone (2018) Figure 5 also demonstrates the tendency of $B_{r}$ and $B_{\phi}$ to change sign at similar heights (our model actually predicts a small offset, due to the viscous contribution to inflow).

### v2

@@ 2.4 Dissipation
Energy is dissipated in our model through both viscosity and magnetic diffusion. The viscous dissipation rate per unit area is simply given by
@@
To calculate the dissipation due to magnetic diffusion, we evaluate
@@
expresses the dissipation of $B_{\phi}$ and the other term refers to $B_{r}$. In discs supported vertically by non-magnetic pressure, as in standard thin discs (with $q\sim O(1)$), $g\gg 1$ and the dissipation due to $B_{\phi}$ dominates. In this limit, with $\xi\sim O(1)$, the magnetic and viscous dissipation rates are comparable. In a steady state, we have
@@
$F_{\rm acc}={1\over 2}\rho rH\bar{v}\Omega^{2}={\xi\over 2}\eta\rho r\Omega^{2}={1\over 8\pi}{GM\dot{M}\over r^{3}}\ ,$ (29)
where $\dot{M}=4\pi\rho rH|\bar{v}|$ is the accretion rate and we have used the accretion speed in the steady-state limit. As noted by Ferreira & Pelletier (1993), the ratio of dissipation to the local release of binding energy is $\sim\xi^{-1}(H/r)$, i.e., small whenever the inflow speed is strongly boosted by magnetic torques.
3 MRI-driven Turbulence
We have seen that the evolution of our simple slab disc model depends critically on the strength of turbulence expressed through the parameter $q$, which represents the ratio of $v_{{\rm A}z}H$ to the turbulent diffusivity $\eta$ or, in physical terms, the ratio of the diffusion time to the Alfvén crossing time. While the complete dependence of the model on $q$ is complex (and probably reflects the model’s limitations), for $q\ga 1$ we see that increasing $q$ (weaker turbulence) leads to stronger advection and retention of magnetic flux, through both faster inflow speeds (due to magnetic extraction of angular momentum) and weaker magnetic diffusion.
Moreover, these behaviors can be very sensitive to the level of turbulence when $q\approx\pi/2$. While the numerical values at which the behaviors set in are undoubtedly an artifact of the simplified model, the qualitative trends are likely robust and can be understood in terms of the slightly different behaviors of mass advection and magnetic flux advection in a diffusive disc, and their delicate balance.
@@ 3 MRI-driven Turbulence
between $\beta_{z}\approx 10^{5}$ and $\beta_{z}\approx 10$, where the parameter $\beta_{z}$ is defined in terms of an imposed vertical magnetic field $B_{z}$ and mid-plane gas pressure through $\beta_{z}=8\pi P_{g}/B_{z0}^{2}$. [^1] This correlation is distinct from the well-known result $\alpha\propto\beta^{-1}$, where in this case $\beta$ is calculated using the magnetic energy density associated with the fully developed turbulence (Blackman et al., 2008). If ${\rm Pr_{t}}$ is a constant, this implies that $\eta\propto v_{{\rm A}z}H$ and $q$ is independent of $\beta_{z}$. Setting $\alpha\sim 10\beta_{z}^{-1/2}$ (Salvesen et al., 2016), this gives $q\sim 0.25{\rm Pr_{t}}$, which falls short of the value $q\approx\pi/2$ needed for a steady state unless Prt is considerably larger than one.
However, the simulations also indicate that the MRI-driven turbulent energy density approaches and exceeds the gas pressure when the vertical magnetic field is still quite weak, at $\beta_{z}\la 10^{2}$. The enhanced turbulent pressure will thicken the disc (Scepi et al., 2024), but also — according to our disc model — drive the Alfvén speed associated with the organized toroidal field, $v_{{\rm A}\phi}$, above the sound speed. In this limit, the growth rate of MRI is suppressed (Blaes & Balbus, 1994). Quasi-linear arguments suggest that the saturated level of MRI-driven turbulence is proportional to the instability growth rate on scales $\sim H$ (Begelman & Armitage, 2023), in which case the amplification of $B_{\phi}$ as $\beta_{z}$ decreases could drive $q$ towards or even beyond $\pi/2$, leading to a steady state (if $q\la\pi/2$) or the rapid accumulation of flux (if $q>\pi/2$). We will adopt the ansatz that for weakly magnetized discs, there is a critical $\beta_{z}$, $\beta_{\rm crit}\sim 10^{2}$, at which $v_{B}=0$ and a steady state can be maintained for both mass flux and magnetic flux.
@@ 4.1 Weakly magnetized discs
Adopting the closure relation for this case, we take $\beta_{z}=\beta_{\rm crit}\equiv 10^{2}\beta_{2}$, implying
@@
This is also very close to the mean accretion speed $|\bar{v}|$, from equation (23). We assume that the disc is supported vertically by gas pressure, implying $H=c_{s}/\Omega$. Interestingly, the reduced level of dissipation inside the disk implies that a magnetically boosted, steady state disc cannot be supported by radiation pressure, as can be seen by the following argument. Vertical support of a disc by radiation pressure requires a local flux $F_{d}(r)=(H/r)GMc/\kappa r^{2}$, where $\kappa$ is the opacity. But we have seen in §2.4 that the actual disc flux is roughly a fraction $H/r$ of the locally liberated binding energy, or $\sim(H/r)GM\dot{M}/8\pi r^{3}$ (cf. equation 29). Comparing the two fluxes, we find that maintaining radiation pressure support requires $\dot{M}$ to be a specific function of radius, $\dot{M}\sim 8\pi cr/\kappa$. This condition clearly cannot be satisfied at all $r$ in a conservative accretion flow, with constant $\dot{M}$. [^2] Expressed as $r\sim\kappa\dot{M}/8\pi c$, this is the condition for the “trapping radius” (Begelman, 1978), where advection and diffusion of radiation are comparable.
The accretion rate can be normalized to the Eddington value using the electron scattering opacity, $\kappa_{\rm es}$:
@@
The dissipation is the sum of contributions from magnetic diffusivity and viscosity, and is given (for $\xi={\rm Pr_{t}}=1$) by
@@
These scalings (and others that can be derived from them) resemble the standard accretion disc scalings in the scattering-dominated regime. In both opacity limits, however, the systematic deviations of the results from standard disc models embody the lower radiative efficiencies of magnetized discs and the suppression of radiation pressure as a major source of vertical support, when $\dot{m}<1$.
@@ 5 Astrophysical Context
A premise of this paper is that accretion discs could be supplied with net magnetic flux from their surroundings. We have argued that the ability of a disk to concentrate flux by advecting it inward is sensitive to its level of magnetization (e.g., through the plasma $\beta$-parameter). In this section we address the global question of whether a disc forced into a highly magnetized state at large radii will then evolve toward a similar state at smaller radii. Given the unique radial dependence of the organized vertical magnetic field, $B_{z}\propto r^{-5/4}$, required for steady accretion in a strongly magnetized disc, this question is clearly crucial for assessing the ability of discs to enter the MAD state.
@@ 5.1 Shakura-Sunyaev–like discs
If $\dot{M}$ is so low that either of the conditions (52) or (53) is satisfied, then the standard Shakura-Sunyaev mode should not be self-consistent because magnetic stresses will take over both the vertical support and angular momentum transport. For accretion rates between the two limiting values, we would expect vigorous turbulence leading to a magnetically supported disc with a magnetically boosted inflow speed, both of which would exacerbate the departure from viscous disk behavior. Paradoxically, in the more extreme limit of condition (53), we face the opposite situation that the level of turbulence might drop significantly as instability is quenched by the strong vertical field; in this case, the disc might effectively stop accreting.
@@ 5.2 Formation of MAD-like states
Like the solution with $\dot{M}_{0}>\dot{M}_{\rm mag}$, this steady flux advection solution also cannot persist to arbitrarily small radii, since the disc scale height satisfies $H/r\propto vr^{1/2}\propto r^{3/4}$. As before, we would expect the strong magnetization assumption to fail when $v\la 0.1c_{s}$, but unlike the case with high $\dot{M}_{0}$, the low-$\dot{M}_{0}$ state cannot revert to a viscous disk solution since strong flux advection continues to be maintained from the outer boundary. Instead, we expect flux to continue to accumulate, working its way inward until the flow approaches a steady state with $\dot{M}\sim\dot{M}_{0}$ and MAD scaling all the way to the central object. Note that the eventual state of the disc would be the true steady state, with $v_{B}\approx 0$, as opposed to the steady flux-advection cases we have been considering.
This picture thus provides a possible scenario for creating a steady-state MAD. However, it assumes that the toroidal field is able to self-regulate to $B_{\phi}\sim v_{K}\rho^{1/2}<B_{z}$. There is no guarantee that this will happen, and we should consider the alternate possibility that the disc “freezes” near $r_{0}$, with so little turbulent activity at these high magnetization levels that $B_{0}$ is inadequate to drive accretion at even the relatively modest rate of $\dot{M}_{0}$. In this case, we would expect the column density near $r_{0}$ to build up until it reaches a level where strong activity resumes, $B_{\phi}$ approaches $\sim 10B_{z}$, and a period of MAD-like accretion with $\dot{M}\sim\dot{M}_{\rm mag}$ occurs. Since the mass supply rate is smaller than $\dot{M}_{\rm mag}$, this would have to be episodic accretion with an on-off duty cycle of $\dot{M}_{0}/\dot{M}_{\rm mag}$.
@@ 6 Discussion and conclusions
We have presented a simple, analytic model for thin accretion discs containing a unidirectional vertical magnetic field, $B_{z}(r)$. Because these discs are threaded by indestructible net magnetic flux, mean-field stochastic dynamo effects are not essential for maintaining the field and we ignore all of the commonly studied terms with the exception of an isotropic turbulent diffusivity, $\eta$, which is related to the turbulent viscosity via a turbulent magnetic Prandtl number, ${\rm Pr}_{t}$. Importantly, we make no claims whatsoever as to the actual role of mean-field or other small-scale dynamo effects; they could well overwhelm the effects we discuss here. However, if they do turn out to be relatively unimportant, our analysis shows that the large-scale field structure can be robustly governed by the balance between field amplification by shear and radial flow, versus turbulent diffusion and dissipation.
To isolate the basic features of the field evolution, we consider a local slab geometry with a scale height $H(r)$, with density $\rho(r)$ and $\eta(r)$ constant for $-H<z<H$ and zero elsewhere. Assuming fixed angular velocity $\Omega(r)$ (taken to be Keplerian in our applications), we calculate the vertical dependences of $B_{r}$, $B_{\phi}$ and the radial velocity explicitly from the induction and angular momentum conservation equations. Both $B_{r}$ and $B_{\phi}$ exhibit sinusoidal behavior about the midplane, with $B_{\phi}$ showing additional linear behavior with $z$ due to the viscous stress and advection of radial magnetic flux.
@@
The sensitivity of the model behavior to $q$ underscores the importance of understanding the relationship between small-scale instabilities and large-scale magnetic structures. Our lack of such understanding is most acute in the strongly magnetized cases. For the level of turbulence we have adopted in this limit — which assumes that the maximum pressure of the organized field is comparable to the turbulent energy density — the squeezing effect of the organized field close to the midplane is of order unity, and is thus quantitatively but not qualitatively important. This begs the question of what happens if $B_{\phi}^{\prime}$ reverses sign within the disc, or indeed how closely the level of turbulence is linked to the local behavior of the organized field. Our closure seems reasonable on energetic grounds, since the same shear flow is responsible for both the turbulent and organized field. But the argument from which the closure relation is derived (Begelman & Armitage, 2023) addresses a more specific context in which the turbulence in a highly magnetized disc is not driven directly by shear via MRI (which Begelman & Armitage 2023 argues becomes too weak when the disc is highly magnetized), but rather by tearing instabilities driven by the vertical gradient of $B_{\phi}$, which in turn has been set up by the shear. Since the strength of the instability depends on $B_{\phi}^{\prime}$, might not the strength of the turbulence and thus $\eta$ be tied to the field structure locally, with dramatic variations as a function of height?
@@ 6.1 Midplane outflows and field reversals
The trigonometric behavior of field strengths and radial velocities with height in our model suggest that discs may exhibit internal radial and toroidal field reversals, accompanied by zones of outward radial flow. Several simulations of magnetized discs have exhibited accretion dominated by regions offset from the midplane (Zhu & Stone, 2018; Lančová et al., 2019; Mishra et al., 2020), accompanied by weak inflow, or even outflow (“backflow”) near the equator. Analyzing the effect in their simulations, Zhu & Stone (2018) argue that this accretion pattern is dominated by the $\phi z$ stress due to large-scale magnetic fields, as earlier noted by Beckwith et al. (2009), rather than by viscosity operating on vertical shear (Kluzniak & Kita, 2000; Guilet & Ogilvie, 2012, 2013). For our magnetically boosted models (i.e. neglecting internal viscosity), we find midplane outflow for $(2n-1)\pi<q<2n\pi$, where $n\geq 1$ is an integer. The corresponding disc exhibits $n$ internal field reversals. Such high values of $q$ would correspond to depressed levels of turbulent diffusivity compared to the fiducial scale $v_{{\rm A}z}H$; it remains to be seen under what conditions this would apply. We note that the source of angular momentum (and energy) driving midplane outflow is the accretion occurring further from the midplane. A portion of the angular momentum liberated in the inflow zone is deposited in the midplane region, driving outflow, while the remainder is carried away by magnetic stresses at the disc surface. In all cases, the vertically integrated mass flux is inward and the layer nearest the disc surface is accreting.
The direction of flow can also be associated with the geometry of the poloidal field. Changing the sign of $v_{r}(0)$ changes the sign of both $B_{r}$ and $B_{\phi}$ near the midplane. Given our magnetic sign convention $B_{z}>0$, equatorial outflow corresponds to $B_{r}<0$, which means that the poloidal field lines projected in the azimuthal plane should bend inward away from the midplane. This is clearly seen in Figure 3 of Mishra et al. (2020), where the field lines bend inward in a region around the midplane characterized by weak outflow. Moreover, our model predicts that $|B_{r}|$ should be maximized where $v_{r}$ changes sign and vanish in the middle of an elevated accretion zone — both effects also apparent in Mishra et al. (2020) Figure 3. This behavior gives the poloidal field lines the appearance of being pinched radially inward at some height above and below the midplane. Analogous information is displayed in Figure 15 of Zhu & Stone (2018) where pinching is also seen, although the relationship between $B_{r}$ and $v_{r}$ is not so clear. Zhu & Stone (2018) Figure 5 also demonstrates the tendency of $B_{r}$ and $B_{\phi}$ to change sign at similar heights (our model actually predicts a small offset, due to the viscous contribution to inflow).